The sum of half Harrys age plus one-third of his age is 50.Find out how old Harry is.
Answer by addingup(3677)

Harry's age: x
1/2x + 1/3x = 50
5/6x = 50
x = 50*6/5
x = 60 years old
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
60 * 1/2 = 30
60 * 1/3 = 20
30 + 20 = 50 Correct